OpenStudy (anonymous):

Hi I have 2 problems that i am stuck with 1) 4 4/5+x= 6 6/15 2) 1 7/8+x= 6 19/24 Please I could really use the help!

OpenStudy (anonymous):

The first problem is 4 4/5+x= 6 11/15

OpenStudy (anonymous):

1. 4 4/5 + x = 6 6/15 24/5 + x = 96/15 x = 96/15 - 24/5 x = 96/15 - 72/15 x = 24/15 = 1 9/15 First I turned the mixed numbers into improper fractions. Then, to get x by itself, I subtracted 24/5 from both sides. And as you should know, when subtracting fractions, you can't subtract unless you have a common denominator. Once you set the equations up to where they have common denominators, you can then subtract. I changed the final answer into a mixed number....However, if your teacher doesn't make you do that, you could leave it as an improper fraction. Now lets see if you can do the 2nd problem...